A pair of vertical angles have degree measure of x2 (x squared) and x + 20. One of the angles can have a measure of ..
a) 20
b) 36
c) 5
d) 25 THE ANSWER IS D)25, but why and how to you get this-please explain in detail
SORRY FOR INCONVENIENCE< BUT YOUR HELP IS MOST APPRECIATED

Answer by fractalier(6550)

Any time straight lines cross, they make congruent vertical angles.
Thus x^2 = x + 20
x^2 - x - 20 = 0
(x - 5)(x + 4) = 0
x = 5 (x cannot be negative)
and if x = 5, the two angles are
x^2 = 25 and
x + 20 = 25